OpenStudy (anonymous):

3sqrt{10}(10sqrt{3}+3sqrt{5})

OpenStudy (espex):

Start by distributing your \(3\sqrt{10}\) in to each term within the parenthesis.

OpenStudy (anonymous):

What's the answer? I just need it desperately right now

OpenStudy (espex):

I can help you through it but I cannot give you the answer, sorry.

OpenStudy (anonymous):

I feel like it's \[13\sqrt{13}+6\sqrt{15}\] but I'm not positive

OpenStudy (mathstudent55):

13sqrt(13) is wrong because you need to multiply 3 and 10 not add them

OpenStudy (mathstudent55):

The same goes for the second product. You need to multiply 3 by 3, and multiply sqrt(3) and sqrt(5). The sqrt part, sqrt(15), is correct, but you need to multiply 3 by 3, not add 3 and 3.

OpenStudy (anonymous):

So is it this? \[30\sqrt{30}+45\sqrt{2}\]

OpenStudy (mathstudent55):